Capital budgeting

The G.Rod Electronic Component is considering replacing a 10 year-old machine that originally cost $37,500, has a current book value of $12,500 with 5 years of expected life left, and is being depreciated using the simplified straight-line method over its 15-year expected life down to a terminal value of zero in 5 years, generating depreciation of $2,500 per year. The replacement machine being considered would cost $100,000 and have a 5-year expected life over which it would be depreciated using the simplified straight-line method down to zero. At termination in 5 years, the new machine would have a salvage value of $35,000. Material efficiencies resulting from the replacement would result in savings of $30,000 per year before depreciation and taxes. Currently, the old machine could be sold for $17,000. Assuming simplified straight-line depreciation, a 34 percent marginal tax rate, and a required rate of return of 20 percent, calculate:

a. The payback period

b. Accounting rate of return

c. The net present value

d. Profitability index

e. Internal rate of return
Answer #1

Year o Year 1 Year 2 Year 3 Year 4 Year 5 | 1 Particulars 2 Capital Outlays 3 Investment 4 Capilal Outluys ($1,00,000,00) {Year 0 Year 1 Year 2 Year 3 Year 4 Year 5 -100000 =SUM(B3:B3) 17000 =-(12500-17000) 348*138 135000 =G11+(5.76%*B3) =44**G12 =

answered by: ANURANJAN SARSAM